OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1 over 2x − 2, find f[g(−6)]. Also can someone explain how to solve the fraction problems so I can try for myself?

OpenStudy (blurbendy):

substitute -6 in for g(x) g(-6) = 1 / ( 2(-6) - 2)

OpenStudy (blurbendy):

= - 1/14

OpenStudy (blurbendy):

then put -1/14 into f(x)

OpenStudy (blurbendy):

f(-1/14) = -1/14 - 4 = -57/14

OpenStudy (anonymous):

9,/ -9/ 3/ -3 Are my answer choices, so do I have to convert this?

OpenStudy (blurbendy):

can you draw the equation for g(x) please

OpenStudy (anonymous):